OpenStudy (anonymous):

what is the cercomfrience of a circle that messers 1.3 feet

OpenStudy (anonymous):

is 1.3 feet the diameter or radius?

OpenStudy (anonymous):

radius

OpenStudy (anonymous):

ok, then circumference is equal to: \[2\pi r\] where r is the radius. so 2*1.3*Pi=2.6*Pi

OpenStudy (anonymous):

thx u

OpenStudy (anonymous):

no problem :)

OpenStudy (amistre64):

gotta teach these kids there manners lol ;)

OpenStudy (anonymous):

amistre are you a math teacher?

OpenStudy (amistre64):

sometimes, when i get bored :)

OpenStudy (anonymous):

that isy father

OpenStudy (amistre64):

i brought the kids along to the library today :)

OpenStudy (anonymous):

good, the library is always nice :) start them on some good math books

OpenStudy (amistre64):

im gonna make them read the discrete math books lol

OpenStudy (anonymous):

haha